# Ziyue Feng ,1,2       Huangying Zhan 11footnotemark: 1 ,1       Zheng Chen 22footnotemark: 2 ,1,3       Qingan Yan 1 Xiangyu Xu 1     Changjiang Cai 1     Bing Li 2     Qilun Zhu 2     Yi Xu 1

### v1

@@ Abstract
We present NARUTO, a neural active reconstruction system that combines a hybrid neural representation with uncertainty learning, enabling high-fidelity surface reconstruction. Our approach leverages a multi-resolution hash-grid as the mapping backbone, chosen for its exceptional convergence speed and capacity to capture high-frequency local features. The centerpiece of our work is the incorporation of an uncertainty learning module that dynamically quantifies reconstruction uncertainty while actively reconstructing the environment. By harnessing learned uncertainty, we propose a novel uncertainty aggregation strategy for goal searching and efficient path planning. Our system autonomously explores by targeting uncertain observations and reconstructs environments with remarkable completeness and fidelity. We also demonstrate the utility of this uncertainty-aware approach by enhancing SOTA neural SLAM systems through an active ray sampling strategy. Extensive evaluations of NARUTO in various environments, using an indoor scene simulator, confirm its superior performance and state-of-the-art status in active reconstruction, as evidenced by its impressive results on benchmark datasets like Replica and MP3D.
@@ 2 Related Work
In autonomous robotics, essential capabilities include localization, mapping, planning, and motion control [56, feng2018model]. These elements have led to research areas like visual odometry [52, 73], stereo matching [62, 20], multi-view stereo [55, 71, 32], structure-from-motion (SfM) [53], and simultaneous localization and mapping [63, 16, 13, 4]. Active SLAM, which combines these approaches for autonomous localization, mapping, and planning, minimizes uncertainties in environmental modeling [12]. We refer readers to the survey papers [4, 33, 45] for a comprehensive discussion regarding the development of active SLAM. Our focus is on active reconstruction, often investigated as exploration problems [64, 17, 3, 34, 58, 39, 57]. a problem that seeks optimal movements for accurate environmental representations [11], primarily for scene and object reconstruction from multiple viewpoints [35, 44, 25, 21, 14, 43].
@@ 3.1 Neural 3D Mapping
[…]
@@ 3.3 Uncertainty-aware Planning
Initially, we set up a multi-level Goal Space, denoted as $\textbf{S}_{g}\in\mathbb{R}^{H\times W\times N}$, comprising layers that are distributed at different heights within the space. The arrangement is such that each layer is approximately 1 meter apart from its adjacent layers, providing a structured vertical distribution throughout the space. Rather than aggregating uncertainties at every vertex within $V_{\sigma}$ onto the Goal Sapce, our method focuses on a set of vertices with the top-k uncertainty, denoted as $\{\textbf{x}_{\sigma}\}^{k}$, where $k=300$. For each point $\textbf{x}_{g}$ sampled within the Goal Space, we accumulate the uncertainty of all visible $\{\textbf{x}_{\sigma}\}^{k}$ points, provided they fall within the optimal observation range of $[0.5,2]m$. Visibility is ascertained by examining the SDF values between $\textbf{x}_{g}$ and $\textbf{x}_{\sigma}$. Upon completing this aggregation process, the goal with the highest aggregated value is subsequently selected as the provisional target location. The goal state $\textbf{s}_{g}$ is defined as the goal location looking at its most uncertain region.
@@ 4.3 Ablation Studies
Figure 6: Replica Results Two scenes (office0, office3) are shown in the first and second rows, respectively. The results represent [ Ground Truth, Uncertainty Net, w/o ActiveRay, Full ]. Our Full method shows better completeness and quality on the highlighted regions.
@@ 7.2 Efficient RRT Details
Furthermore, E-RRT enhances the process of node expansion. Instead of adding a single node, our method integrates a series of feasible points uniformly distributed between a randomly generated node and its nearest neighbor in the tree, based on a predefined step size, for instance, 10 cm. This modification substantially accelerates the expansion of the tree, especially in the initial growth stages.
@@ 8 Runtime Analysis
In this section, we present a detailed runtime analysis of the three major modules in NARUTO, as illustrated in Fig. 8. The first module is a simulator for data generation. The second is a mapping module optimized for a hybrid scene representation. Lastly, we have an uncertainty-aware planning module. For data generation, HabitatSim requires, on average, 24.4ms to generate $680\times 1200$ RGB-D data per iteration. The Mapping module, although taking about 300ms per iteration, averages 60.5ms since it is activated only every five keyframes. The Active Planning module averages 2.1ms, which includes 0.3ms for collision detection per iteration. Additionally, Active Planning encompasses two modules that are triggered occasionally when the ’PLAN_REQUIRED’ condition is met. These are the uncertainty-aware goal searching, averaging 6.8ms, and RRT path planning, averaging 8.1ms. In conclusion, our analysis demonstrates that NARUTO offers real-time capabilities, particularly due to its efficient planning module.
@@ 9.1 Detailed results on MP3D and Replica
Method Metric Gdvg gZ6f HxpK pLe4 YmJk Avg.
ANM [70] MAD (cm) $\downarrow$ 3.77 3.18 7.03 3.25 4.22 4.29
Acc. (cm) $\downarrow$ 5.09 4.15 15.60 5.56 8.61 7.80
Comp. (cm) $\downarrow$ 5.69 7.43 15.96 8.03 8.46 9.11
Comp. Ratio $\uparrow$ 80.99 80.68 48.34 76.41 79.35 73.15
Ours MAD (cm) $\downarrow$ 1.60 1.23 1.53 1.37 1.45 1.44
Acc. (cm) $\downarrow$ 3.78 3.36 9.24 5.15 10.04 6.31
Comp. (cm) $\downarrow$ 2.91 2.31 2.67 3.24 3.86 3.00
Comp. Ratio $\uparrow$ 91.15 95.63 91.62 87.76 84.74 90.18

### v2

# NARUTO: Neural Active Reconstruction from Uncertain Target Observations

@@ Abstract
We present NARUTO, a neural active reconstruction system that combines a hybrid neural representation with uncertainty learning, enabling high-fidelity surface reconstruction. Our approach leverages a multi-resolution hash-grid as the mapping backbone, chosen for its exceptional convergence speed and capacity to capture high-frequency local features. The centerpiece of our work is the incorporation of an uncertainty learning module that dynamically quantifies reconstruction uncertainty while actively reconstructing the environment. By harnessing learned uncertainty, we propose a novel uncertainty aggregation strategy for goal searching and efficient path planning. Our system autonomously explores by targeting uncertain observations and reconstructs environments with remarkable completeness and fidelity. We also demonstrate the utility of this uncertainty-aware approach by enhancing SOTA neural SLAM systems through an active ray sampling strategy. Extensive evaluations of NARUTO in various environments, using an indoor scene simulator, confirm its superior performance and state-of-the-art status in active reconstruction, as evidenced by its impressive results on benchmark datasets like Replica and MP3D. Project page: oppo-us-research.github.io/NARUTO-website/
@@ 2 Related Work
In autonomous robotics, essential capabilities include localization, mapping, planning, and motion control [64]. These elements have led to research areas like visual odometry [60, 83], monocular depth estimation [20, 82, 3, 24, 23], multi-view stereo [70, 28, 63, 80, 40, 6, 11], structure-from-motion (SfM) [61], path planning [27, 35, 34, 22], and SLAM [71, 19, 16, 5, 75]. Active SLAM, which combines these approaches for autonomous localization, mapping, and planning, minimizes uncertainties in environmental modeling [15]. We refer readers to the survey papers [5, 41, 53] for a comprehensive discussion regarding the development of active SLAM. Our focus is on active reconstruction, often investigated as exploration problems [72, 21, 4, 42, 66, 47, 65]. a problem that seeks optimal movements for accurate environmental representations [14], primarily for scene and object reconstruction from multiple viewpoints [43, 52, 33, 29, 17, 51].
@@ 3.1 Neural 3D Mapping
where $w_{i}=\varphi(\frac{s_{i}}{tr})\varphi(-\frac{s_{i}}{tr})$ represents the weights computed along the ray, obtained by applying Sigmoid functions $\varphi(.)$ to the predicted SDF $\textit{s}_{i}$ within a truncated range $tr=10\text{cm}$.
@@ 3.3 Uncertainty-aware Planning
Initially, we set up a multi-level Goal Space, denoted as $\textbf{S}_{g}\in\mathbb{R}^{H\times W\times N}$, comprising layers that are distributed at different heights within the space. The arrangement is such that each layer is approximately 1 meter apart from its adjacent layers, providing a structured vertical distribution throughout the space. Rather than aggregating uncertainties at every vertex within $V_{\sigma}$ onto the Goal Space, our method focuses on a set of vertices with the top-k uncertainty, denoted as $\{\textbf{x}_{\sigma}\}^{k}$, where $k=300$. For each point $\textbf{x}_{g}$ sampled within the Goal Space, we accumulate the uncertainty of all visible $\{\textbf{x}_{\sigma}\}^{k}$ points, provided they fall within the optimal observation range of $[0.5,2]m$. Visibility is ascertained by examining the SDF values between $\textbf{x}_{g}$ and $\textbf{x}_{\sigma}$. Upon completing this aggregation process, the goal with the highest aggregated value is subsequently selected as the provisional target location. The goal state $\textbf{s}_{g}$ is defined as the goal location looking at its most uncertain region.
@@ 4.3 Ablation Studies
Figure 6: Replica Results Two scenes (office0, office3) are shown in the first and second rows, respectively. The results represent [ Ground Truth, Uncertainty Net, w/o ActiveRay, Full ]. Our Full method shows a better completeness and quality on the highlighted regions. Note that the GT visualization uses view-dependent rendering, unlike our mapping backbone, resulting in color differences in the visualizations.
@@ 7 Implementation Details
Memory requirement
Memory consumption varies depending on the scene size. As a reference, in a $120m^{3}$ scene, the corresponding GPU memory and RAM are 8.1GB and 8.6GB respectively. The consumption can be further reduced with a more efficient implementation as our current implementation involves intensive exchanges between RAM and GPU memories.
@@ 7.2 Efficient RRT Details
Furthermore, E-RRT enhances the process of node expansion. Instead of adding a single node, our method integrates a series of feasible points uniformly distributed between a randomly generated node and its nearest neighbor in the tree, based on a predefined step size, for instance, 10 cm, up the distance of $M\times\text{step size}$. Here $M$ equals to 10. This modification substantially accelerates the expansion of the tree, especially in the initial growth stages.
@@ 8 Runtime Analysis
8.1 System Runtime
In this section, we present a detailed runtime analysis of the three major modules in NARUTO, as illustrated in Fig. 8. The first module is a simulator for data generation. The second is a mapping module optimized for a hybrid scene representation. Lastly, we have an uncertainty-aware planning module. For data generation, HabitatSim requires, on average, 24.4ms to generate $680\times 1200$ RGB-D data per iteration. The Mapping module, although taking about 300ms per iteration, averages 60.5ms since it is activated only every five keyframes. The Active Planning module averages 2.1ms, which includes 0.3ms for collision detection per iteration. Additionally, Active Planning encompasses two modules that are triggered occasionally when the ’PLAN_REQUIRED’ condition is met. These are the uncertainty-aware goal searching, averaging 6.8ms, and RRT path planning, averaging 5.77 ms. In conclusion, our analysis demonstrates that NARUTO offers real-time capabilities, particularly due to its efficient planning module.
8.2 RRT Runtime Analysis
Method Time (ms) Node Num. Step Num.
RRT $19\times 10^{3}$ $19\times 10^{3}$ $28\times 10^{3}$
w/o direct line $17\times 10^{3}$ $20\times 10^{3}$ $21\times 10^{3}$
w/o fast tree $16.00$ $44.17$ $2.56$
Ours (E-RRT) $5.77$ $16.70$ $1.19$
Table 3: RRT runtime analysis on Replica-room0. We conducted a runtime analysis of RRT variants, revealing that our optimized RRT implementation significantly outpaces traditional RRT in planning speed, achieving real-time planning capabilities.
In this section, we delve deeper into our optimized RRT implementation, as outlined in Sec. 7.2. We have engineered a customized version of RRT that enhances planning speed through several strategies:
  • Direct Line: Actively identifying straight paths that link the RRT tree to the goal.
  • Fast Tree: Speeding up the expansion of the tree.
  • Parallel Computing: Utilizing GPU processing for increased efficiency.
These innovations significantly reduce the time required for path planning, making our RRT variant highly suitable for real-time applications. We present an ablation study on the runtime performance of our RRT approach in Tab. 3. To maintain consistency, all experiments were conducted using parallel processing for nearest-neighbor searches during tree expansion.
Evaluation
Our evaluation of the methods encompasses three key metrics: the average time taken for each path planning request, the average number of nodes generated within the RRT tree, and the average number of steps taken in the RRT process.
Analysis
Compared to traditional RRT, our efficient RRT implementation is markedly faster, both in average planning time and iteration count. It also generates fewer nodes and uses less memory, as shown by the reduced average number of nodes required per planning request. The ablation study detailed in Tab. 3 highlights that our primary strategy for improvement involves identifying potential straight paths, drawing inspiration from RRT-Connect [34]. This approach, along with quicker tree growth, not only accelerates the planning process but also decreases memory usage.
@@ 9.1 Detailed results on MP3D and Replica
Method Metric Gdvg gZ6f HxpK pLe4 YmJk Avg.
ANM [79] MAD (cm) $\downarrow$ 3.77 3.18 7.03 3.25 4.22 4.29
Acc. (cm) $\downarrow$ 5.09 4.15 15.60 5.56 8.61 7.80
Comp. (cm) $\downarrow$ 5.69 7.43 15.96 8.03 8.46 9.11
Comp. Ratio $\uparrow$ 80.99 80.68 48.34 76.41 79.35 73.15
Ours MAD (cm) $\downarrow$ 1.60 1.23 1.53 1.37 1.45 1.44
Acc. (cm) $\downarrow$ 3.78 3.36 9.24 5.15 10.04 6.31
Comp. (cm) $\downarrow$ 2.91 2.31 2.67 3.24 3.86 3.00
Comp. Ratio $\uparrow$ 91.15 95.63 91.62 87.76 84.74 90.18
@@
Method Metrics office0 office1 office2 office3 office4 room0 room1 room2
CoSLAM Comp. Ratio $\uparrow$ 96.33 94.65 92.47 91.43 91.34 94.67 95.45 93.95
(no tracking) Traj. (m) $\uparrow$ 18.20 11.56 23.16 29.16 25.22 24.69 16.21 23.07
Ours Comp. Ratio $\uparrow$ 98.17 95.26 97.54 93.91 97.93 98.28 98.04 98.47
Traj. (m) $\uparrow$ 81.27 30.02 90.20 88.59 96.36 73.91 96.99 41.31
Table 6: Per-scene trajectory length evaluation on Replica[67] dataset
@@
9.2 More qualitative comparison on MP3D
Figure 24: More Matterport3D results We trim the reconstruction results for a better comparison. Compared to the baseline method, ANM [79], our method shows more precise and complete reconstructions.
For the completeness of the study, we provide more comparison between ground truth, ANM baseline [79], and our method in Matterport3D dataset, as shown in Fig. 24. We trim the meshes for a better visualization purpose.
9.3 Comparison against passive mapping methods
Figure 25: Qualitative comparison between active and passive mapping methods. For Co-SLAM [73], we disable the tracking thread and run the reconstruction using a pre-defined trajectory. Active NARUTO shows a more complete and precise reconstruction, especially for the regions that have not been adequately covered by the passive scanning.
In traditional mapping methods, typically involving environments scanned by human-operated sensing devices, the trajectory of scanning significantly impacts the reconstruction’s quality and completeness. Such approaches are termed passive mapping methods, characterized by the absence of a planning or guidance module. In Tab. 2, we present a quantitative comparison between Passive Neural Mapping and Active Neural Mapping, utilizing Co-SLAM as the backbone. Here, we aim to offer additional qualitative comparisons in Fig. 25 to highlight differences in reconstruction details more vividly. In passive Co-SLAM (with tracking disabled), regions may be missed or poorly reconstructed if not adequately covered by the scanning trajectory. Conversely, our active reconstruction method ensures a more comprehensive and accurate reconstruction, effectively addressing these limitations.
We compared the trajectory lengths of passive versus active scanning on the Replica dataset, with the results detailed in Tab. 6. Under the same conditions (2000 frames with 400 keyframes), passive scanning may result in redundant observations due to the lack of guided scanning. Active scanning, on the other hand, enables more extensive coverage and yields superior reconstruction quality. However, this approach typically results in longer trajectories, as the agent continuously moves to ensure comprehensive scanning of the environment.